# MIDAS: Multi-level Intent, Domain, And Slot Knowledge Distillation for Multi-turn NLU

### v2

@@ Abstract
Although Large Language Models(LLMs) can generate coherent and contextually relevant text, they often struggle to recognise the intent behind the human user’s query. Natural Language Understanding (NLU) models, however, interpret the purpose and key information of user’s input to enable responsive interactions. Existing NLU models generally map individual utterances to a dual-level semantic frame, involving sentence-level intent and word-level slot labels. However, real-life conversations primarily consist of multi-turn conversations, involving the interpretation of complex and extended dialogues. Researchers encounter challenges addressing all facets of multi-turn dialogue conversations using a unified single NLU model. This paper introduces a novel approach, MIDAS, leveraging a multi-level intent, domain, and slot knowledge distillation for multi-turn NLU. To achieve this, we construct distinct teachers for varying levels of conversation knowledge, namely, sentence-level intent detection, word-level slot filling, and conversation-level domain classification. These teachers are then fine-tuned to acquire specific knowledge of their designated levels. A multi-teacher loss is proposed to facilitate the combination of these multi-level teachers, guiding a student model in multi-turn dialogue tasks. The experimental results demonstrate the efficacy of our model in improving the overall multi-turn conversation understanding, showcasing the potential for advancements in NLU models through the incorporation of multi-level dialogue knowledge distillation techniques.
@@ 1 Introduction
Natural Language Understanding (NLU) within the realm of Natural Language Processing (NLP) explores the mechanisms through which computers comprehend human language. Developing a hierarchical semantic framework encompassing domain, intent, and slot has become pivotal in representing the meaning embedded in natural language [1]. We present a conversation example that shows the way of annotation for word-level slots, sentence-level intent, and conversation-level domain from the M2M dataset in Figure 1-(a). The dialogue consists of a total of 9 turns, and each turn includes word-level slot tokens and sentence-level intent information, and the dialogue corresponds to one domain, ‘restaurant’.
Large Language Models (LLM) have received lots of attention in generating human-like text based on user prompts. However, they are still limited when it comes to deeper communication and diverse key information [^1] We tested NLU benchmarks with several LLMs, including LLaMa2, Gemma, and GPT3.5, visualised in Appendix G.1 and G.2. Hence, we investigate how to improve the state-of-the-art existing NLU techniques. While existing NLU literature predominantly concentrates on single-turn utterances within a single domain, recent advancements in multi-turn datasets have paved the way for annotations at the dialogue level, spanning across diverse domains. Interpreting more extended and intricate conversations with multiple turns necessitates understanding the ongoing context and retaining previously gathered information. Traditional NLU involves mapping single utterances to a dual-level semantic structure, encompassing sentence-level intent and word-level slot labels. With real-life conversations extending across multiple turns, there is an evident demand for research incorporating dialogue history, as demonstrated by improved performance through dialogue context. The challenge extends beyond dual-level understanding to encompass a three-level comprehension: sentence-level intent, word-level slot, and conversation-level domain classification. However, researchers encounter challenges in handling all aspects of multi-turn dialogue conversations through a single unified NLU model, due to computational complexity and a lack of distillability of multi-level knowledge.
(a)                                                                           (b)
Figure 1: (a) An example of conversations with word-level slots, sentence-level intents, and conversation-level domain annotation from M2M. B-NP (B-Number of People), B-RN (B-Restaurant Name), O (Others) (b) The multi-level teacher knowledge distillation framework for the multi-turn NLU task. Note that we applied three multi-level teachers: Intent Detection, Slot Filling, and Domain Classification. In this framework, we conduct diverse Loss objectives, including $\mathcal{L}_{rel}$, $\mathcal{L}_{sim}$, $\mathcal{L}_{sce}$, $\mathcal{L}_{tp}$ and $\mathcal{L}_{KD}$, which represent relation loss, similarity loss, student cross-entropy loss, teacher prediction supervise loss, and Kullback-Leibler Divergence loss, respectively.
Model   Year   Token   Sentence   Document Dialogue Type   Joint Integration
@@
Table 1: Summary of existing joint NLU models and ours. Token, Sentence, and Document columns indicate whether the relevant information is used for joint integration. KD refers to knowledge distillation. The complete set of summary tables is detailed in Appendix A
Addressing this need, our paper introduces a novel multi-level multi-teacher knowledge distillation model to enhance NLU understanding in multi-turn dialogues, leveraging diverse levels of knowledge embedded in these datasets. Notably, our model stands as the pioneering approach in multi-teacher knowledge distillation, catering to distinct facets of knowledge within a dialogue. To achieve this, our approach involves the construction of teachers at different levels, specifically focusing on sentence-level intent detection, word-level slot filling, and conversation-level domain classification. We fine-tune these multi-level teachers to acquire the relevant knowledge and combine these to educate the student model in dialogue tasks facilitated by novel multi-level teacher loss functions.
There are two major contributions: First, we introduce a novel multi-level, multi-teacher knowledge distillation model to enhance multi-turn NLU. It outperforms across widely-used multi-NLU datasets, producing superior performance in all intent detection, slot filling, and domain classification, even compared with the LLMs. Secondly, we introduce multi-level teacher loss functions, shedding light on their impact within the multi-teacher knowledge distillation and guiding a student model.
@@ 2 Related works
There is a large body of NLU modelling literature, and we briefly introduce the joint NLU models and knowledge distillation models in this section. A more detailed summary of joint NLU models and our model can be found in Table 1.
Natural Language Understanding Early investigations aimed to tackle various Natural Language Understanding tasks by developing models for slot filling and intent detection separately. It has now become commonplace to enhance joint models using transfer learning [10, 11]. One notable strategy involves fine-tuning a language model [12, 13, 14] to address the limited generalisation capability caused by insufficient data and leverage high-quality representations from the language model. In these studies, the majority classify intent through the [CLS] representation and slots through each token representation [15, 16]. Another approach to transfer learning involves training a model through knowledge distillation, where a smaller model, known as the student model, is trained to mimic the behaviour of a larger, more complex model, referred to as the teacher model. The predominant method is self-distillation, where both the teacher and student models share the same structures [7, 17]. However, those models mainly focused on single-turn dialogue or adopted only one teacher model. As real-life conversations involve multiple turns, there is a growing need for research that incorporates dialogue history. In multi-turn dialogues, it has been demonstrated that encoding the dialogue history enhances performance [3, 9, 18, 19]. Hence, our model is the first method for multi-teacher knowledge distillation to teach different aspects of knowledge in dialogue. As NLU consists of multiple tasks such as intent classification and slot filling, it is more appropriate to train the student model using a specialised teacher for each task. To the best of our knowledge, there have been no attempts to employ multi-teacher distillation for multi-turn-based NLU.
@@
Knowledge Distillation Knowledge Distillation (KD) defines a learning approach involving using a well-trained network of teachers to guide the training of a student network for various tasks. Early KD transfers knowledge from one teacher to one student model [20]. Multi-teacher KD, inspired by ensemble learning, aims to enhance performance by incorporating knowledge from multiple teacher models into a student model [21, 22, 23, 24, 25, 26, 27, 28]. It was common for KD to use multiple teachers to learn the same domain, regardless of whether the teacher had the same or a different architecture. Recent methods have been proposed in which each teacher learns a different domain and imparts knowledge to the student [29, 30]. Additionally, methods for learning different modalities have also been proposed, which are classified into two types: the teacher and student learn different modalities, a concept known as cross-modality [31, 32] and the teacher learns different modalities, and the student receives all modalities [33].
3 MIDAS: Multi-level intent, domain, and slot knowledge distillation for multi-turn NLU
@@
We propose a new multi-level dialogue teacher knowledge distillation framework, MIDAS, that trains the student model $S$ with multi-level teachers to enhance the Natural Language Understanding (NLU) capabilities. Note that we have three multi-level dialogue knowledge teachers, including intent detection, slot filling and domain topic classification. To achieve this, we initially construct teachers with distinct levels of dialogue knowledge, denoted as $T=\{T_{ID},T_{SF},T_{DC}\}$, where $T$ is the set of teacher models, and $ID$, $SF$, and $DC$ correspond to Intent Detection, Slot Filling, Domain Classification tasks. Then, we fine-tune the teacher models $T$ to acquire knowledge from each task. Finally, a combination of all three multi-level teachers $T$ is employed to instruct the student model $S$ in dialogue tasks using our newly proposed multi-teacher knowledge loss objectives. The comprehensive architecture is depicted in Figure 1-(b).
@@ 3.1 Multi-Level teacher construction
We first construct the teachers of different dialogue document component understanding levels, including word-level slot, sentence-level intent, and conversation-level domain knowledge. The inputs for all teachers consist of utterances from each turn in dialogue datasets, denoted by $X^{i}={x^{i}_{1},x^{i}_{2},...,x^{i}_{l}}$, where $X^{i}$ represents the $i_{th}$ utterance in the entire dataset, $l$ is the length of the utterance, and $x^{i}_{l}$ signifies a word in the utterance.
@@
Word-level slot filling teacher $T_{SF}$ predicts the slot type for each word, providing knowledge to the student model about key slots in the dialogue. The output of $T_{SF}$ is denoted as $\hat{Y}^{i}_{SF}={\hat{Y}^{i}_{SF,1},\hat{Y}^{i}_{SF,2},...,\hat{Y}^{i}_{SF,l}}$, representing the predicted slot types for each word, where $\hat{Y}^{i}_{SF,l}\in{0,1,...,k_{SF}-1}$, and $k_{SF}$ is the number of slot types.
@@
Sentence-level intent detection teacher $T_{ID}$ predicts the intent of the utterance, aiding the student model in comprehending the overall intent of each turn. The prediction of $T_{ID}$ is symbolised as $\hat{Y}^{i}_{ID}$, where $\hat{Y}^{i}_{ID}\in{0,1,...,k_{ID}-1}$, and $k_{ID}$ represents the number of intents in the dataset.
@@
Conversation document-level domain classification teacher $T_{DC}$ forecasts the dialogue’s domain, providing knowledge to classify it and understand its background knowledge. The prediction from $T_{DC}$ is indicated as $\hat{Y}^{i}_{DC}$, where $\hat{Y}^{i}_{DC}\in{0,1,...,k_{DC}-1}$, and $k_{DC}$ denotes the number of domains in the dataset.
@@
Using these three levels of teachers, our objective is to instruct the student model to comprehend dialogues from multiple perspectives, incorporating word-level slot knowledge, sentence-level intent, and document-level domain background knowledge. By doing so, we enhance the student model’s grasp of dialogues across various levels. There are two primary reasons for utilising multi-level dialogue knowledge teachers to train a student. First, individually deploying a pre-trained model for each task consumes more computational resources, and some machines may not support running multiple pre-trained models. Instead, the knowledge distillation process leads to more robust models and is resistant to adversarial attacks. Incorporating soft targets from the teacher model can help the student model learn smoother decision boundaries. Secondly, we posit that diverse levels of knowledge derived from multi-turn conversation understanding datasets can enhance the comprehension of each specific natural language understanding task, surpassing the benefits of learning from single-level dialogue knowledge. Note that we use pre-trained models as the foundational structure for our teachers. After experimenting with various backbones, we determined that BERT yields one of the best results overall, as detailed in Section 5.2. These pre-trained models undergo fine-tuning using specific data for each level, resulting in distinct teachers with expertise in intent detection, slot filling, and document classification. Pre-trained models, having been trained on extensive text data, exhibit the capacity to transfer knowledge effectively. Ultimately, we leverage the collective knowledge of these refined teachers to train the student model comprehensively.
@@ 3.2 Multi-level Teacher Fine-tuning
We perform separate fine-tuning of pre-trained models on intent detection (ID), slot-filling (SF), and domain classification (DC) tasks. This yields multi-level teachers, $T_{ID}$, $T_{SF}$, and $T_{DC}$ respectively, corresponding to sentence-level, word-level and sentence-level knowledge, respectively. Each pre-trained model specialises in learning knowledge at one specific level from the dialogue datasets, resulting in teachers possessing different levels of dialogue document component understanding. It’s important to note that each teacher focuses on one level of dialogue knowledge. This approach is motivated by two factors. First, learning knowledge from a single task is less complex than incorporating knowledge from all tasks, simplifying the fine-tuning of pre-trained models. Secondly, instead of burdening a single model with the challenge of mastering knowledge from all aspects of dialogues, each teacher focuses on a specific level of understanding, such as word-level slot filling, sentence-level intent detection, or document-level domain classification. For each task, we consolidate data from two datasets (MultiWOZ and M2M) by merging split and corresponding label sets. For example, the training set for fine-tuning includes data from both datasets. We apply cross-entropy loss and fine-tune the pre-trained models for a fixed number of epochs, utilising the checkpoint from the last epoch as the teacher model. The process is described as follows:
$X_{j,train}$ $=$ $X^{1}_{j,train},...,X^{N_{M2M}}_{j,train},X^{1}_{j,train},...,X^{N_{MultiWoz}}_{j,train},$
$\mathcal{L}_{tce}$ $=$ $cross\_entropy(T_{j}(X_{j,train}),Y_{j}),$
$j$$\in$ $\{DC,ID,SF\},$

### v3

@@ Abstract
Although Large Language Models (LLMs) can generate coherent text, they often struggle to recognise user intent behind queries. In contrast, Natural Language Understanding (NLU) models interpret the purpose and key information of user input for responsive interactions. Existing NLU models typically map utterances to a dual-level semantic frame, involving sentence-level intent (SI) and word-level slot (WS) labels. However, real-life conversations primarily consist of multi-turn dialogues, requiring the interpretation of complex and extended exchanges. Researchers encounter challenges in addressing all facets of multi-turn dialogue using a unified NLU model. This paper introduces MIDAS, a novel approach leveraging multi-level intent, domain, and slot knowledge distillation for multi-turn NLU. We construct distinct teachers for SI detection, WS filling, and conversation-level domain (CD) classification, each fine-tuned for specific knowledge. A multi-teacher loss is proposed to facilitate the integration of these teachers, guiding a student model in multi-turn dialogue tasks. Results demonstrate the efficacy of our model in improving multi-turn conversation understanding, showcasing the potential for advancements in NLU through multi-level dialogue knowledge distillation. Our implementation is open-sourced on GitHub [^1] https://github.com/adlnlp/Midas.
MIDAS: Multi-level Intent, Domain, And Slot
Knowledge Distillation for Multi-turn NLU
Yan Li1,3, So-Eon Kim2, Seong-Bae Park2, Soyeon Caren Han1,3,thanks: Corresponding Author, 1The University of Sydney, 2Kyung Hee University, 3The University of Melbourne 1yali3816@uni.sydney.edu.au, 2{sekim0211, sbpark71}@khu.ac.kr, 3caren.han@unimelb.edu.au
@@ 1 Introduction
Natural Language Understanding (NLU) within the realm of Natural Language Processing (NLP) explores the mechanisms through which computers comprehend human language. Developing a hierarchical semantic framework encompassing domain, intent, and slot has become pivotal in representing the meaning embedded in natural language Weld et al. (2022). We present a conversation example that shows the way of annotation for WSs, SI, and CD from the M2M dataset in Figure 1. The dialogue consists of a total of 9 turns, and each turn includes WS tokens and SI information, and the dialogue corresponds to one domain, ‘restaurant’.
Figure 1: An example of conversations with WSs, SIs, and CD annotation from M2M. B-NP (B-Number of People), B-RN (B-Restaurant Name), O (Others).
Model   Year   Word   Sentence   Document Dialogue Type   Joint Integration
@@
[…]
Table 1: Summary of existing joint NLU models and ours. Word, Sentence, and Document columns indicate whether the relevant information is used for joint integration. KD refers to knowledge distillation. The complete set of summary tables is detailed in Appendix A.
Large Language Models (LLM) have received much attention in generating human-like text based on user prompts. However, they are still limited when it comes to deeper communication and diverse key information [^2] We tested NLU benchmarks with several LLMs, including LLaMa2, LLama3.1, Gemma, QWen2, GPT3.5, and GPT4o visualised in Appendix D.2.1 and D.2.2.. Hence, we investigate how to improve the state-of-the-art existing NLU techniques. While existing NLU literature predominantly concentrates on single-turn utterances within a single domain, recent advancements in multi-turn datasets have paved the way for annotations at the dialogue level, spanning across diverse domains. Interpreting more extended and intricate conversations with multiple turns necessitates understanding the ongoing context and retaining previously gathered information. Traditional NLU involves mapping single utterances to a dual-level semantic structure, encompassing SI and WS labels. With real-life conversations extending across multiple turns, there is an evident demand for research incorporating dialogue history, as demonstrated by improved performance through dialogue context. The challenge extends beyond dual-level understanding to encompass a three-level comprehension: SI, WS, and CD classification. However, researchers encounter challenges in handling all aspects of multi-turn dialogue conversations through a single unified NLU model, due to computational complexity and a lack of distillability of multi-level knowledge.
This paper introduces a novel multi-level multi-teacher knowledge distillation model to enhance NLU understanding in multi-turn dialogues, leveraging diverse levels of knowledge embedded in these datasets. Notably, our model is the pioneering approach in multi-teacher knowledge distillation, catering to distinct facets of knowledge within a dialogue. To achieve this, our approach involves the construction of teachers at different levels, specifically focusing on SI detection, WS filling, and CD classification. We fine-tune these multi-level teachers to acquire the relevant knowledge and combine these to educate the student model in dialogue tasks facilitated by novel multi-level teacher loss functions. There are two major contributions:
1) We introduce a novel multi-level, multi-teacher knowledge distillation model to enhance multi-turn NLU. It outperforms widely-used multi-NLU datasets, producing superior performance in all intent detection, slot filling, and domain classification, even compared with the LLMs.
2) We introduce multi-level teacher loss functions, shedding light on their impact within the multi-teacher knowledge distillation and guiding a student model.
@@ 2 Related works
There is a large body of NLU modelling literature, and we briefly introduce the joint NLU models and knowledge distillation models. A summary of these models and our model are in Table 1.
Natural Language Understanding Early works addressed slot filling and intent detection separately. Current research commonly employs joint models with transfer learning Rongali et al. (2021); M’hamdi et al. (2021), where fine-tuning language models Dao et al. (2021); Abro et al. (2022); Mei et al. (2023) enhances generalisation by leveraging high-quality representations. Typically, intent is classified through the [CLS] token and slots through individual token embeddings Chen et al. (2019); Han et al. (2021a); Heo et al. (2022); Luo and Feng (2024). Another transfer learning strategy is knowledge distillation, where a smaller student model learns from a larger teacher model, often using self-distillation Chen et al. (2022b); Cheng et al. (2023). However, these methods primarily address single-turn dialogues or use only one teacher model. Multi-turn dialogues benefit from encoding dialogue history, leading to performance gains Bapna et al. (2017); Weld et al. (2023); Wu and Juang (2023); Tu et al. (2023); Chen et al. (2024). Our model is the first to employ multi-teacher knowledge distillation for multi-turn NLU. Distinct teachers specialise in intent classification, slot filling, and domain classification, thus effectively distilling multi-level knowledge. To the best of our knowledge, no previous work has explored multi-teacher distillation in multi-turn dialogue NLU.
@@
Knowledge Distillation Knowledge Distillation (KD) defines a framework where a well-trained teacher network guides the training of a student network for various tasks. Traditional KD employs a single teacher to train one student model Hinton et al. (2014). Multi-teacher KD, inspired by ensemble learning, integrates knowledge from multiple teachers to enhance student model performance Wu et al. (2021a); Yuan et al. (2021); Jung et al. (2023); Wang et al. (2021b); Huang et al. (2023); Amirkhani et al. (2021); Mirzadeh et al. (2020); Son et al. (2021). Typically, multiple teachers focus on the same domain, regardless of architecture. Recent approaches involve each teacher specialising in different domains and imparting domain-specific knowledge to the student Pan et al. (2021); Ji et al. (2023). Cross-modality methods have been explored: either the teacher and student learn different modalities Kong et al. (2019); Ni et al. (2022), or the teacher learns multiple modalities, and the student receives all modalities Jin et al. (2021).
Figure 2: The proposed multi-level teacher knowledge distillation framework for the multi-turn NLU task. Note that we applied three multi-level teachers: Intent Detection, Slot Filling, and Domain Classification. In this framework, we conduct diverse Loss objectives, including $\mathcal{L}_{rel}$, $\mathcal{L}_{sim}$, $\mathcal{L}_{sce}$, $\mathcal{L}_{tp}$ and $\mathcal{L}_{KD}$, which represent relation loss, similarity loss, student cross-entropy loss, teacher prediction supervise loss, and Kullback-Leibler Divergence loss, respectively.
3 MIDAS
@@
We propose a new multi-level dialogue teacher knowledge distillation framework, MIDAS, that trains the student model $S$ with multi-level teachers to enhance the NLU capabilities. We have three multi-level dialogue knowledge teachers, including intent detection, slot filling and domain topic classification. To achieve this, we initially construct teachers with distinct levels of dialogue knowledge, denoted as $T=\{T_{ID},T_{SF},T_{DC}\}$, where $T$ is the set of teacher models, and $ID$, $SF$, and $DC$ correspond to Intent Detection, Slot Filling, Domain Classification. Then, we fine-tune the teacher models $T$ to acquire knowledge from each task. On top of simple KD losses, we also introduce two novel loss functions, relation loss $\mathcal{L}_{rel}$ and teacher prediction supervised loss $\mathcal{L}_{tp}$, specifically designed for this domain. These facilitate knowledge transfer from multi-level teachers to the student model. The overall architecture is illustrated in Figure 2.
@@ 3.1 Multi-Level teacher construction
We first construct the teachers of different dialogue document component understanding levels, including WS, SI, and CD knowledge. The inputs for all teachers consist of utterances from each turn in dialogue datasets, denoted by $X^{i}={x^{i}_{1},x^{i}_{2},...,x^{i}_{l}}$, where $X^{i}$ represents the $i_{th}$ utterance in the entire dataset, $l$ is the length of the utterance, and $x^{i}_{l}$ signifies a word in the utterance.
@@
1) Word-level teacher $T_{SF}$ predicts the slot type for each word, providing knowledge to the student model about key slots in the dialogue. The output of $T_{SF}$ is $\hat{Y}^{i}_{SF}={\hat{Y}^{i}_{SF,1},\hat{Y}^{i}_{SF,2},...,\hat{Y}^{i}_{SF,l}}$, representing the predicted slot types for each word, where $\hat{Y}^{i}_{SF,l}\in{0,1,...,k_{SF}-1}$, and $k_{SF}$ is the number of slot types.
@@
2) Sentence-level teacher $T_{ID}$ predicts the intent of the utterance, aiding the student model in comprehending the overall intent of each turn. The prediction of $T_{ID}$ is symbolised as $\hat{Y}^{i}_{ID}$, where $\hat{Y}^{i}_{ID}\in{0,1,...,k_{ID}-1}$, and $k_{ID}$ represents the number of intents in the dataset.
@@
3) Conversation Document-level teacher $T_{DC}$ forecasts the dialogue’s domain, providing knowledge to classify it and understand its background knowledge. The prediction of $T_{DC}$ is indicated as $\hat{Y}^{i}_{DC}$, where $\hat{Y}^{i}_{DC}\in{0,1,...,k_{DC}-1}$, and $k_{DC}$ denotes the number of domains in the dataset.
@@
Using these three levels of teachers, our objective is to instruct the student model to comprehend dialogues from multiple perspectives, incorporating WS, SI, and CD background knowledge. By doing so, we enhance the student model’s grasp of dialogues across various levels. There are two primary reasons for utilising multi-level dialogue knowledge teachers to train a student. First, individually deploying a pre-trained model for each task consumes more computational resources, and some machines may not support running multiple pre-trained models. Instead, the knowledge distillation process leads to more robust models and is resistant to adversarial attacks. Incorporating soft targets from the teacher model can help the student model learn smoother decision boundaries. Secondly, we posit that diverse levels of knowledge derived from multi-turn conversation understanding datasets can enhance the comprehension of each specific natural language understanding task, surpassing the benefits of learning from single-level dialogue knowledge. Note that we use pre-trained models as the foundational structure for our teachers. After experimenting with various backbones, we determined that BERT yields one of the best results overall, as detailed in Section 5.2. These pre-trained models undergo fine-tuning using specific data for each level, resulting in distinct teachers with expertise in intent detection, slot filling, and document classification. Pre-trained models, having been trained on extensive text data, exhibit the capacity to transfer knowledge effectively. Ultimately, we leverage the collective knowledge of these refined teachers to train the student model comprehensively.
@@ 3.2 Multi-level Teacher Fine-tuning
We perform separate fine-tuning of pre-trained models on $ID$, $SF$, and $DC$ tasks. This yields multi-level teachers, $T_{ID}$, $T_{SF}$, and $T_{DC}$ respectively, corresponding to sentence-level, word-level and sentence-level knowledge, respectively. Each pre-trained model specialises in learning knowledge at one specific level from the dialogue datasets, resulting in teachers possessing different levels of dialogue document component understanding. It’s important to note that each teacher focuses on one level of dialogue knowledge. This approach is motivated by two factors. First, learning knowledge from a single task is less complex than incorporating knowledge from all tasks, simplifying the fine-tuning of pre-trained models. Secondly, instead of burdening a single model with the challenge of mastering knowledge from all aspects of dialogues, each teacher focuses on a specific level of understanding, such as WS filling, SI detection, or CD classification. For each task, we consolidate data from two datasets (MultiWOZ and M2M) by merging split and corresponding label sets. For example, the training set for fine-tuning includes data from both datasets. We apply cross-entropy loss and fine-tune the pre-trained models for a fixed number of epochs, utilising the checkpoint from the last epoch as the teacher model. The process is described as follows:
$\begin{split}X_{j,tr}=&X^{1}_{j,tr},\dots,X^{N_{M2M}}_{j,tr},X^{1}_{j,tr},\dots,\\ &X^{N_{MWOZ}}_{j,tr},\\ \mathcal{L}_{tce}=&cross\_entropy(T_{j}(X_{j,tr}),Y_{j}),\\ j\in&\{DC,ID,SF\}\end{split}$